OpenStudy (anonymous):

simplify 1. b8/b1 2. (3.8x10^5)divided by (1.9x10^-2) 3.(4.5x10^6)divided by (1.6x10^3) 4. 4^9/4^6 i really need help on this

OpenStudy (anonymous):

note that when dividing similiar basis - subtract powers or exponents so for b^8 / b^1 8-1= 7 answer is b^7

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

so what did you get for the 2nd one?

OpenStudy (anonymous):

i dont know how to do it can you explain ? for number 4 i got 4^4

OpenStudy (anonymous):

actually 9-6 is 3 so you should have gotton 4^3

OpenStudy (anonymous):

for the 2nd one: 3.8 / 1.9 = (use calc if you can not see this as 2) 5 - (-2) = 5 + 2 = 7 answer is 2 x 10^7 you follow? ok you try #3